# SAM-guided Graph Cut for 3D Instance Segmentation

### v2

@@ Abstract
This paper addresses the challenge of 3D instance segmentation by simultaneously leveraging 3D geometric and multi-view image information. Many previous works have applied deep learning techniques to 3D point clouds for instance segmentation. However, these methods often failed to generalize to various types of scenes due to the scarcity and low-diversity of labeled 3D point cloud data. Some recent works have attempted to lift 2D instance segmentations to 3D within a bottom-up framework. The inconsistency in 2D instance segmentations among views can substantially degrade the performance of 3D segmentation. In this work, we introduce a novel 3D-to-2D query framework to effectively exploit 2D segmentation models for 3D instance segmentation. Specifically, we pre-segment the scene into several superpoints in 3D, formulating the task into a graph cut problem. The superpoint graph is constructed based on 2D segmentation models, where node features are obtained from multi-view image features and edge weights are computed based on multi-view segmentation results, enabling the better generalization ability. To process the graph, we train a graph neural network using pseudo 3D labels from 2D segmentation models. Experimental results on the ScanNet, ScanNet++ and KITTI-360 datasets demonstrate that our method achieves robust segmentation performance and can generalize across different types of scenes. Our project page is available at https://zju3dv.github.io/sam_graph.
$*{}^{*}$qual contribution. ${}^{\dagger}$ork done during internship at Zhejiang University.
@@ 1 Introduction
Figure 1: Comparison of the generalization ability. By constructing superpoint graph with SAM [25], our 3D segmentation method demonstrates commendable generalizability. After training on ScanNet, not only do we achieve good performance on its validation set, but we can also effectively generalize to data collected with different devices (ScanNet++) and even to entirely different types of scenes (KITTI-360).
Instance segmentation of 3D scenes is a cornerstone of many applications, such as augmented and virtual reality, robot navigation, and autonomous driving. One typical pipeline for 3D scene segmentation is using deep neural networks[37, 38, 29, 48] to process the point cloud of the target scene to predict segmentation results. These methods[28, 11, 52, 56, 19, 12, 23, 50, 42] generally require annotated point clouds for training. However, annotating point clouds is costly, and there is a lack of datasets with the large scale and diversity similar to 2D image datasets. As a result, these methods are often limited to specific types of scenes and struggle to generalize to in-the-wild scenes.
@@
Recently, some 2D-to-3D lifting methods [57, 53, 54, 26, 44, 1, 58] have used a bottom-up framework, which first runs 2D segmentation on each view to obtain several masks, and then attempts to establish correspondences among masks across different views. These masks are then merged in 3D to obtain the 3D segmentation results. However, such bottom-up approach has a significant issue: 2D segmentation masks from different views may be inconsistent. For example, some instances may be segmented in some views but missing in others, thus severely degrading the performance of 3D segmentation.
In this paper, we propose a novel 3D segmentation approach based on a 3D-to-2D query framework, which effectively utilizes 2D segmentation models. Unlike previous methods that generate multiple masks from multi-view images first, our approach begins with a pre-segmentation of the 3D scene into several superpoints. Then, we construct a superpoint graph of the target scene and transform the problem into graph segmentation. The edge weights of the graph are obtained by projecting graph nodes onto multiple views, using the prompt mechanism of SAM [25] to predict multi-view masks, and calculating the intersection of corresponding masks. The node features are obtained by aggregating multi-view image features. Finally, we use a graph neural network to regress the edge affinity for the graph partition [46]. The SAM-based graph enables better generalization ability, as shown in Fig. 1. In addition, we develop a scheme to generate pseudo 3D labels from a 2D segmentation network and design a training strategy to effectively leverages these pseudo labels, allowing us to train our model without any manual 3D annotations.
We conduct experiments on the ScanNet dataset, training the graph neural network on the training set and evaluating on the validation set, achieving state-of-the-art results. Without any fine-tuning, we also evaluate our method on ScanNet++ and KITTI-360 datasets. The experimental results indicate that, with the guidance of SAM in the construction of the graph, our method demonstrates excellent generalization capabilities. It can generalize across different types of data, encompassing different data acquisition methods and scene types.
@@
  • We propose a novel 3D-to-2D-query framework that leverages SAM to construct node features and edge weights of the superpoint graph, effectively improving the generalization ability of the graph segmentation.
  • We develop a scheme to generate pseudo 3D labels from a 2D segmentation network, enabling our model to be trained without any manual 3D annotations.
  • We achieve robust segmentation results on ScanNet [6] and effectively generalize to ScanNet++ [59] and KITTI-360 [30] datasets.
@@ 2 Related work
The goal of 3D scene segmentation is to group the scene point cloud into semantically meaningful regions or distinct objects. Previous works leverage large-scale 3D labeled datasets to accomplish this objective in a supervised manner. They first train a neural network to extract per-point features, and then assign one predicted label to each point based on the extracted features. [16, 32, 21, 51, 47, 38, 29, 55, 48] achieve semantic segmentation on point cloud, and [28, 11, 52, 56, 19, 12, 23, 50] further distinguish between different objects with the same semantics, thus getting 3D instance segmentation results. Recently, Mask3D[42] leverage Transformer[49] to construct the segmentation network, attaining superior instance segmentation quality on 3D point clouds. 3D-SIS[20] performs 3D instance segmentation on RGB-D scan data, fusing image features extracted from 2D convolution networks with 3D scan geometry features, allowing accurate inference for object bounding boxes, class labels, and instance masks.
@@
Some works exploit 2D large vision-language models to achieve open-vocabulary segmentation in 3D space. OpenScene[36] back-project per-pixel image features extracted by large vision language models from multi-view posed images to form a feature point cloud endowed with open-vocabularity abilities for various down-stream scene understanding tasks. PLA[10] constructs multi-scale 3D-text pairs and uses contracstive learning to enable the model to learn language-aware embeddings for 3D semantic and instance segmentation.
@@
Partitioning 3D point cloud into a collection of small, geometrically homogeneous regions, coined superpoints, can yield a decent rough prediction and effectively simplify the process of scene segmentation. [45] and [22] propose to represent each 3D scene by constructing a superpoint graph, where superpoints serve as graph nodes, and then based on this graph, 3D instance segmentation is performed by learning inter-superpoint affinity and clustering superpoint nodes into 3D objects.
Figure 2: Overview of our pipeline. Our proposed 3D instance segmentation pipeline consists of three main parts. 1. We over-segment the input mesh / point cloud into superpoints and construct the structure of the superpoint graph based adjacency (Sec. 3.1). 2. We rely on the prompt mechanism of SAM [25] to annotate the nodes and edges of the graph (Sec. 3.2). The multi-view SAM backbone features corresponding to each superpoint are aggregated into node features. The intersection ratio between the multi-view SAM masks corresponding to each pairwise superpoints constituting an edge is used to annotate the weight of the edge. The higher the intersection ratio, the higher the edge weight. 3. We use a graph neural network to further process the SAM-annotated graph and perform graph cut based on the calculated edge affinity scores to obtain the instance segmentation results (Sec. 3.3).
2D-to-3D lifting.
@@
Due to the lower cost of acquiring and annotating 2D images, the scale and diversity of 2D annotated datasets [9, 31, 27, 5, 13, 18], are much larger compared to 3D datasets, facilitating the emergence of many highly effective 2D segmentation methods [25, 24, 39, 4] in recent years, making the use of 2D segmentation for 3D tasks a new and promising approach. Semantic-NeRF [61], based on the NeRF framework, utilizes outputs from a 2D semantic segmentation network at each view to train a 3D semantic field. Since it can fuse information from multiple viewpoints, this method is robust to inaccuracies and noise in individual view segmentations, yielding better 3D semantic segmentation results. However, extending this approach directly to instance segmentation is challenging, which is more complex since instance IDs in multi-view image segmentation results may not be consistent, making it necessary to design an effective label matching mechanisms in order to lift multi-view 2D instance segmentation results to 3D space.
@@
To address this issue, [44] solves a linear assignment for instance identifiers across views with machine generated semantic and instance labels as supervision, [1] proposes a scalable slow-fast clustering objective function to fuse 2D predictions into a unified 3D scene segmentation results represented with a neural field, but they fail to produce satisfactory results for complex scenes with a large number of instances. SAD [2] uses SAM to segment both images and depth maps, combining the advantages of both for improved results. SAM3D [58] proposes a method for point cloud fusion. It segments each frame and gradually merges the segmentation results of all frames together. For scenes with known geometry, it can achieve segmentation very efficiently. Some works integrate SAM [25] into the Neural Radiance Fields (NeRF) [33] framework for 3D segmentation. For example, [3] merges SAM features into Instant NGP [34] in 3D, allowing users to segment an object from 3D space through multiple clicks. OR-NeRF [60] enables users to segment an object by clicking and then remove it from the scene.
@@
Additionally, some methods combine segmentation and reconstruction, enabling separate reconstruction of each object. [57] propose to decompose a scene by learning an object-compositional nerual radiance field, with each standalone object separated from the scene and encoded with a learnable object activation codes, allowing more flexible downstream applications. To cope with the ambiguity of conventional volume rendering pipelines, [53, 54] further utilizes the Signed Distance Function (SDF) representation to exert explicit surface constraint.[26] represents each object in the scene with a small MLP and builds an object-level dense SLAM that detects objects on-th-fly and dynamically adds them to its map.
@@ 3 Method
Given 3D geometry and calibrated multi-view images of a scene, our goal is to obtain the 3D instance segmentation of the scene. In this paper, we propose a novel segmentation framework, as illustrated in Fig. 2. We first perform over-segmentation on the 3D geometry to generate a set of superpoints, reformulating the task into a graph cut problem (Sec. 3.1). Then, Sec. 3.2 describes how to leverage SAM to construct node features and edge weights of the superpoint graph. In Sec. 3.3, we introduce a graph neural network for 3D segmentation, which is trained with pseudo labels generated by 2D segmentation predictions.
@@ 3.1 Building the superpoint graph
In an indoor or outdoor scene, not only can we easily acquire multi-view images, but we can also obtain the scene’s geometry (either in point clouds or mesh form) through depth cameras/laser scanners (for indoor scenes) or lidar (for large-scale outdoor scenes). With the geometry of the scene available, we can proceed to pre-segment the scene based on traditional methods to obtain a set of superpoints. For mesh, we apply the method in [6], which calculating the similarity between mesh vertices based on their normal directions and then conduct the graph cut algorithm [14]. For point clouds, we employ the method in [17], which first compute a local geometric feature vector (dimensionality and verticality) for each point, then perform Potts energy segmentation [8].
@@ 3.2 Constructing edge weights and node features
To accomplish segmentation of the 3D scene, our primary task is to determine whether two superpoints on each edge in graph should be merged. To this end, we leverage multi-view image information and employ SAM to annotate the graph so that we can send the annotated graph into neural network for segmentation. Specifically, we utilize the prompt mechanism of SAM to annotate the edges and use SAM encoder features to annotate the nodes.
Prompt mechanism of SAM.
Unlike previous 2D instance segmentation methods that take an image as input and output its segmentation map, SAM (Segment Anything Model) [25] operates by taking an image and a prompt as inputs and producing corresponding segmentation result. A typical prompt could be one or several 2D points. The image and the prompt are fed into an image encoder and a prompt encoder separately. Subsequently, a transformer-based decoder computes the cross-attention between prompt features and image features to generate the mask. Specifically, SAM can output multiple valid masks with associated confidence scores. In our experiments, we tend to prefer masks with a larger area because they are more likely to represent an entire object. We only resort to selecting masks with a relatively smaller area when the confidence of the larger mask is low (please refer to the supplementary materials for detailed implementation).
Figure 3: Relationship of coefficient and 2D superpoints distance. For two superpoints, their distance in 2D images will be farther under near and frontal views than faraway and collinear views. We assume that SAM achieves better performance on near and frontal views. Thus, we consider the 2D distance as a factor in calculating the coefficient of each view.
Edge weights.
The prompt mechanism of SAM introduces flexibility to calculate the edge weights between two 3D superpoints. Part 2.2 of Fig. 2 presents an illustration of computing the weight. Specifically, we first select a view where both superpoints are visible (if such a view does not exist, we regard the weight to be zero). Then, the two superpoints are projected onto the image space and $k$ points are uniformly sampled in each projection to serve as prompts for running SAM, thereby obtaining a mask for each superpoint. For masks corresponding to the two superpoints, we examine their intersection situation. Specifically, if two superpoints of an edge correspond to masks $A$ and $B$, then we calculate the edge weight as $w=\max(\frac{A\bigcap B}{A},\frac{A\bigcap B}{B})$, where $A\bigcap B$ denotes the intersection of $A$ and $B$.
@@
Node features.
In addition to edge weights, we also annotate node features with SAM. Specifically, for each node in the graph model, we identify all views that observe the corresponding superpoint. Within the projection range of that superpoint in each view, we randomly sample several points, interpolate to obtain features extracted by the SAM encoder and average the features obtained from all views to represent the attributes of that vertex. Part 2.1 of Fig. 2 presents an example of obtaining node features.

### v3

@@ Abstract
Equal contribution Corresponding authors
This paper addresses the challenge of 3D instance segmentation by simultaneously leveraging 3D geometric and multi-view image information. Many previous works have applied deep learning techniques to 3D point clouds for instance segmentation. However, these methods often failed to generalize to various types of scenes due to the scarcity and low-diversity of labeled 3D point cloud data. Some recent works have attempted to lift 2D instance segmentations to 3D within a bottom-up framework. The inconsistency in 2D instance segmentations among views can substantially degrade the performance of 3D segmentation. In this work, we introduce a novel 3D-to-2D query framework to effectively exploit 2D segmentation models for 3D instance segmentation. Specifically, we pre-segment the scene into several superpoints in 3D, and formulate the task into a graph cut problem. The superpoint graph is constructed based on 2D segmentation models, enabling great segmentation performance on various types of scenes. We employ a GNN to further improve the robustness, which can be trained using pseudo 3D labels generated from 2D segmentation models. Experimental results on the ScanNet200, ScanNet++ and KITTI-360 datasets demonstrate that our method achieves state-of-the-art segmentation performance. Our project page is available at https://zju3dv.github.io/sam_graph.
Keywords:
3D Instance Segmentation 3D Scene Understanding Graph Neural Network
@@ 1 Introduction
Figure 1: Thanks to the inductive bias introduced by SAM-annotated superpoint graph, our method achieves good segmentation performance and generalization capabilities. After training solely on ScanNet200, our model can effectively generalize to data collected with different devices (ScanNet++) and even to entirely different types of scenes (KITTI-360).
Instance segmentation of 3D scenes is a cornerstone of many applications, such as augmented and virtual reality, robot navigation, and autonomous driving. One typical pipeline for 3D scene segmentation is using deep neural networks[39, 40, 31, 54] to process the point cloud of the target scene to predict segmentation results. These methods[30, 13, 58, 64, 21, 14, 25, 56, 48] generally require annotated point clouds for training. However, annotating point clouds is costly, and thus there is a lack of datasets with the large scale and diversity similar to 2D image datasets. As a result, these methods are often limited to specific types of scenes and struggle to generalize to in-the-wild scenes.
@@
Recently, some 2D-to-3D lifting methods [63, 59, 60, 28, 50, 3, 65] have used a bottom-up framework, which first runs 2D segmentation on each view to obtain several masks, and then attempts to establish correspondences among masks across different views. These masks are then merged in 3D to obtain the 3D segmentation results. However, such a bottom-up approach has a significant issue: 2D segmentation masks from different views may be inconsistent. For example, some instances may be segmented in some views but missing in others, thus severely degrading the performance of 3D segmentation.
In this paper, we propose a novel 3D segmentation approach based on a 3D-to-2D query framework, which effectively utilizes 2D segmentation models. Unlike previous methods that generate multiple masks from multi-view images first, our approach begins with a pre-segmentation of the 3D scene into several superpoints. Then, we construct a superpoint graph of the target scene and transform the problem into graph cut. The edge weights of the graph are obtained by projecting graph nodes onto multiple views, using the prompt mechanism of SAM [27] to predict multi-view masks, and calculating the intersection of corresponding masks. The node features are obtained by aggregating multi-view image features. Finally, we use a graph neural network to refine the edge affinity for the graph partition [52]. The SAM-based graph enables great segmentation performance on various types of scenes. In addition, we develop a scheme to generate pseudo 3D labels from a 2D segmentation network and design a training strategy to effectively leverages these pseudo labels, allowing us to train our model without any manual 3D annotations.
We conduct experiments on the ScanNet200, ScanNet++ and KITTI-360 datasets. The experimental results indicate that, with the guidance of SAM in the construction of the graph, our method achieves good segmentation results on various types of scenes. Moreover, the GNN module in our method has good generalization capabilities, which is able to generalize well to other datasets with significant differences after trained on one dataset with pseudo-labels.
@@
  • We propose a novel 3D-to-2D-query framework that leverages SAM to construct node features and edge weights of the superpoint graph, enabling great segmentation performance on various types of scenes.
  • We employ GNN to improve the robustness and develop a scheme to generate pseudo 3D labels from a 2D segmentation network, enabling our model to be trained without any manual 3D annotations.
  • We demonstrate state-of-the-art performance on ScanNet200 [45], ScanNet++ [66] and KITTI-360 [32] datasets.
@@ 2 Related work
3D scene segmentation.  The goal of 3D scene segmentation is to group the scene point cloud into semantically meaningful regions or distinct objects. Previous works leverage large-scale 3D labeled datasets to accomplish this objective in a supervised manner. They first train a neural network to extract per-point features, and then assign one predicted label to each point based on the extracted features. [18, 34, 23, 57, 53, 40, 31, 61, 54] achieve semantic segmentation on point cloud, and [30, 13, 58, 64, 21, 14, 25, 56] further distinguish between different objects with the same semantics, thus getting 3D instance segmentation results. Recently, Mask3D[48] leverage Transformer[55] to construct the segmentation network, attaining superior instance segmentation quality on 3D point clouds. 3D-SIS[22] performs 3D instance segmentation on RGB-D scan data, fusing image features extracted from 2D convolution networks with 3D scan geometry features, allowing accurate inference for object bounding boxes, class labels, and instance masks.
@@
Some works exploit 2D large vision-language models to achieve open-vocabulary segmentation in 3D space. OpenScene[38] back-project per-pixel image features extracted by large vision language models from multi-view posed images to form a feature point cloud endowed with open-vocabularity abilities for various down-stream scene understanding tasks. PLA[12] constructs multi-scale 3D-text pairs and uses contrastive learning to enable the model to learn language-aware embeddings for 3D semantic and instance segmentation.
@@
Partitioning 3D point cloud into a collection of small, geometrically homogeneous regions, coined superpoints, can yield a decent rough prediction and effectively simplify the process of scene segmentation. [51] and [24] propose to represent each 3D scene by constructing a superpoint graph, where superpoints serve as graph nodes, and then based on this graph, 3D instance segmentation is performed by learning inter-superpoint affinity and clustering superpoint nodes into 3D objects. [43] propose an efficient semantic segmentation for large-scale scenes by partitioning point clouds into a hierarchical superpoint structure, and [44] further extend it for panoptic segmentation by scalable graph clustering.
@@
2D-to-3D lifting.  Due to the lower cost of acquiring and annotating 2D images, the scale and diversity of 2D annotated datasets [11, 33, 29, 7, 15, 20], are much larger compared to 3D datasets, facilitating the emergence of many highly effective 2D segmentation methods [27, 26, 41, 6] in recent years, making the use of 2D segmentation for 3D tasks a new and promising approach. Semantic-NeRF [69], based on the NeRF framework, utilizes outputs from a 2D semantic segmentation network at each view to train a 3D semantic field. Since it can fuse information from multiple viewpoints, this method is robust to inaccuracies and noise in individual view segmentations, yielding better 3D semantic segmentation results. However, extending this approach directly to instance segmentation is challenging, which is more complex since instance IDs in multi-view image segmentation results could inconsistent, making it necessary to design an effective label matching mechanisms in order to lift multi-view 2D instance segmentation results to 3D space.
@@
To address this issue, [50] solves a linear assignment for instance identifiers across views with machine generated semantic and instance labels as supervision, [3] proposes a scalable slow-fast clustering objective function to fuse 2D predictions into a unified 3D scene segmentation results represented with a neural field. SAD [4] uses SAM to segment both images and depth maps, combining the advantages of both for improved results. SAM3D [65] proposes a method for point cloud fusion. It segments each frame and gradually merges the segmentation results of all frames together. For scenes with known geometry, it can achieve segmentation very efficiently. SAI3D [67] partitions a 3D scene into geometric primitives, which are then progressively merged into 3D instance segmentations that are consistent with the multi-view SAM masks. MaskClustering [62] proposes a novel metric called view consensus to better exploit multi-view observation for 3D instance segmentation. Some works integrate SAM [27] into the Neural Radiance Fields (NeRF) [35] framework for 3D segmentation. For example, [5] merges SAM features into Instant NGP [36] in 3D, allowing users to segment an object from 3D space through multiple clicks. OR-NeRF [68] enables users to segment an object by clicking and then remove it from the scene.
@@
Additionally, some methods combine segmentation and reconstruction, enabling separate reconstruction of each object. [63] propose to decompose a scene by learning an object-compositional neural radiance field, with each standalone object separated from the scene and encoded with a learnable object activation codes, allowing more flexible downstream applications. To cope with the ambiguity of conventional volume rendering pipelines, [59, 60] further utilizes the Signed Distance Function (SDF) representation to exert explicit surface constraint.[28] represents each object in the scene with a small MLP and builds an object-level dense SLAM that detects objects on-the-fly and dynamically adds them to its map.
@@ 3 Method
Figure 2: Overview of our pipeline. Our proposed 3D instance segmentation pipeline consists of three main parts. 1. We over-segment the input mesh / point cloud into superpoints and construct the structure of the superpoint graph based on adjacency (Sec. 3.1). 2. We utilize the prompt mechanism of SAM [27] to annotate the nodes and edges of the graph (Sec. 3.2). The node features are aggregated from multi-view SAM backbone features corresponding to each superpoint. The edge weights are calculated based on the intersection ratio between the multi-view SAM masks corresponding to each pair of superpoints that constitute an edge. 3. We use a graph neural network to further process the SAM-annotated graph and perform graph cut based on the calculated edge affinity scores to obtain the instance segmentation results (Sec. 3.3).
Given 3D geometry and calibrated multi-view images of a scene, our goal is to obtain its 3D instance segmentation. In this paper, we propose a novel segmentation framework, as illustrated in Fig. 2. We first perform over-segmentation on the 3D geometry to generate a set of superpoints, reformulating the task into a graph cut problem (Sec. 3.1). Then, Sec. 3.2 describes how to leverage SAM to construct node features and edge weights of the superpoint graph. In Sec. 3.3, we introduce a graph neural network for 3D segmentation, which is trained with pseudo labels generated by 2D segmentation predictions.
@@ 3.1 Building the superpoint graph
In an indoor or outdoor scene, not only can we easily acquire multi-view images, but we can also obtain the scene’s geometry (either in point clouds or mesh form) through depth cameras/laser scanners (for indoor scenes) or LiDAR (for large-scale outdoor scenes). With the geometry of the scene available, we can proceed to pre-segment the scene based on traditional methods to obtain a set of superpoints. For mesh, we apply the method in [8], which calculates the similarity between mesh vertices based on their normal directions and then conducts the graph cut algorithm [16]. For point clouds, we employ the method in [19], which first computes a local geometric feature vector (dimensionality and verticality) for each point, then performs Potts energy segmentation [10].
@@ 3.2 Constructing edge weights and node features
To accomplish segmentation of the 3D scene, our primary task is to determine whether two superpoints on each edge should be merged. To this end, we leverage multi-view image information and employ SAM to annotate the graph so that we can apply graph cut for segmentation. Specifically, we utilize the prompt mechanism of SAM to annotate the edges and use SAM encoder features to annotate the nodes.
Prompt mechanism of SAM.  Unlike previous 2D instance segmentation methods that take an image as input and output its segmentation map, SAM (Segment Anything Model) [27] operates by taking an image and a prompt as inputs and producing corresponding segmentation result. A typical prompt could be one or several 2D points. The image and the prompt are fed into an image encoder and a prompt encoder separately. Subsequently, a transformer-based decoder computes the cross-attention between prompt features and image features to generate the mask. Specifically, SAM can output multiple valid masks with associated confidence scores. In our experiments, we tend to prefer masks with a larger area because they are more likely to represent an entire object. We only resort to selecting masks with a relatively smaller area when the confidence of the larger mask is low (please refer to the supplementary materials for detailed implementation).
Figure 3: Relationship of coefficient and 2D superpoints distance. For two superpoints, their distance in 2D images will be farther under near and frontal views than faraway or collinear views. We assume that SAM achieves better performance on near and frontal views. Thus, we consider the 2D distance as a factor in calculating the coefficient of each view.
Edge weights.  The prompt mechanism of SAM introduces flexibility to calculate the edge weights between two 3D superpoints. Part 2.2 of Fig. 2 presents an illustration of computing the weight. Specifically, we first select a view where both superpoints are visible (if such a view does not exist, we regard the weight to be zero). Then, the two superpoints are projected onto the image space and $k$ points are uniformly sampled in each projection to serve as prompts for running SAM, thereby obtaining a mask for each superpoint. For masks corresponding to the two superpoints, we examine their intersection situation. Specifically, if two superpoints of an edge correspond to masks $A$ and $B$, then we calculate the edge weight as $w=\max(\frac{A\bigcap B}{A},\frac{A\bigcap B}{B})$, where $A\bigcap B$ denotes the intersection of $A$ and $B$.
@@
Node features.  In addition to edge weights, we also annotate node features with SAM. Specifically, for each node in the graph model, we identify all views that observe the corresponding superpoint. Within the projection range of that superpoint in each view, we randomly sample several points, interpolate to obtain features extracted by the SAM encoder and average the features obtained from all views to represent the attributes of the node. Part 2.1 of Fig. 2 presents an example of obtaining node features.